OpenStudy (anonymous):

help please show steps. Thanks -4(x^-3y^-1)-2

OpenStudy (king):

can u post it using equation editor.....?

OpenStudy (anonymous):

i just tried. I dont see how u put exponents in the editor/

OpenStudy (king):

\[8(x ^{-3y ^{-1}})\]

OpenStudy (king):

is this rite....

OpenStudy (anonymous):

almost. the is needs to be a -4 and the and at the end it need to be ^-2

jhonyy9 (jhonyy9):

-4(1/(yx^3))-2 = -4/(yx^3) -2 = -(4+2yx^3)/(yx^3)

jhonyy9 (jhonyy9):

meljaytunalan (-2) from the end is exponent ?